### v2

# Transformer-based RGB-T Tracking with Channel and Spatial Feature Fusion

Abstract
Complementary RGB and TIR modalities enable RGB-T tracking to achieve competitive performance in challenging scenarios. Therefore, how to better fuse cross-modal features is the core issue of RGB-T tracking. Some previous methods either insufficiently fuse RGB and TIR features, or depend on intermediaries containing information from both modalities to achieve cross-modal information interaction. The former does not fully exploit the potential of using only RGB and TIR information of the template or search region for channel and spatial feature fusion, and the latter lacks direct interaction between the template and search area, which limits the model’s ability to fully exploit the original semantic information of both modalities. To alleviate these limitations, we explore how to improve the performance of a visual Transformer by using direct fusion of cross-modal channels and spatial features, and propose CSTNet. CSTNet uses ViT as a backbone and inserts cross-modal channel feature fusion modules (CFM) and cross-modal spatial feature fusion modules (SFM) for direct interaction between RGB and TIR features. The CFM performs parallel joint channel enhancement and joint multilevel spatial feature modeling of RGB and TIR features and sums the features, and then globally integrates the sum feature with the original features. The SFM uses cross-attention to model the spatial relationship of cross-modal features and then introduces a convolutional feedforward network for joint spatial and channel integration of multimodal features. Comprehensive experiments show that CSTNet achieves state-of-the-art performance on three public RGB-T tracking benchmarks. To promote the practicality of the model, we retrain the model with CSNet as the pre-training weights in the model with CFM and SFM removed, and propose CSTNet-small, which achieves 36% reduction in parameters and 24% reduction in Flops, and 50% speedup with a 1-2% performance decrease. Code is available at https://github.com/LiYunfengLYF/CSTNet.
Index Terms:
RGB-T tracking, Channel Feature Fusion, Spatial Feature Fusion, Transformer Network.
I Introduction
Figure 1: Comparison between our cross-modal feature fusion method with other Transformer-based RGB-T trackers. (a) TIR features are fused into RGB features by extracting prompts, such as ViPT [1] (b) RGB and TIR features interact through the intermediate tokens, such as TBSI [2] (c) Our method achieves cross-modal interaction through channel and spatial feature fusion of RGB and TIR features.
RGB-T tracking is a combination of single object tracking (SOT) and multi-modal vision task that provides multi-modal appearance templates of the target in the initial frame and then predicts the position and scale of the target in the multi-modal images in subsequent frames. It combines the advantages of the visual modality, which contains more information, and the infrared modality, which is less affected by the environment, to achieve stable target tracking in low light and adverse weather conditions. Due to its advantage in processing visual and infrared image sequences, it has a wide range of potential applications in traffic monitoring [3], robot perception [4] and autonomous driving [5].
As a multi-modal vision task, how to achieve effective cross-modal interaction plays a crucial role in RGB-T tracking. It is worth noting that while some trackers [6] discuss the impact of RGB-T image misalignment on tracking performance, most current trackers are designed with the assumption of spatial registration of RGB-T images. In previous works, MDNet-based trackers have proposed a large number of cross-modal feature fusion methods, such as channel feature fusion [7], adapter [8], challenge-aware model [9], spatio-temporal feature fusion [10], multi-scale feature fusion [11], etc. These methods greatly advance the development of RGB-T tracking. However, the MDNet architecture itself makes it difficult for these trackers to extract discriminative deep features. Then, Siamese-based trackers [12][13][14] further explore the channel fusion methods of cross-modal deep features. However, their sample candidate boxes (RoIs) lack a global receptive field , because only local multimodal features of the search area are used for fusion. Overall, these MDNet-based and Siamese-based methods are limited by their own architecture and do not fully explore the potential of channel and spatial feature fusion. With the introduction of Transformer in RGB-T tracking, some methods use visual prompt learning [1][15] or adapter [16] to achieve better modeling of RGB feature guided by TIR feature, as illustrated in Figure 1 (a), while others are similar to the above methods, as illustrated in Figure 1 (b), like using feature bridging fusion [2], spatio-temporal feature fusion [17], etc. to achieve effectively model of cross-modal features. Although these methods effectively achieve cross-modal feature interaction based on transformers, they are implemented by extracting intermediate prompt features or using partially fused cross-modal tokens. Their indirect feature interaction leads to a lack of use of the original semantics of multi-modal features.
Based on the above analysis, although cross-modal channel feature fusion has been introduced in some MDNet-based and Siamese-based trackers, the multimodal features they extract limit the performance of the fusion module. For Transformer-based methods, although the multimodal features become better, channel and spatial feature fusion based on direct feature interaction does not receive sufficient attention.
As a basic feature fusion method, channel and feature fusion has significant advantages in the pixel-level interaction between RGB and TIR features. It directly models the original semantics of RGB and TIR features, thus avoiding the loss of key features caused by the use of intermediate tokens. For other RGB-T visual tasks, such as RGB-T semantic segmentation and RGB-T saliency detection, many cross-modal channel and spatial feature fusion modules have been proposed for Transformer-based methods, such as using cross-attention to model cross-modal feature representations [18], using channel attention and spatial feature fusion module to fuse intermediate features from different stages [19], etc. The success of the above feature fusion module demonstrates the enormous potential of direct interaction between RGB and TIR features in Transformer.
Therefore, this paper aims to combine channel and feature fusion with Visual Transformer (ViT). Specifically, we investigate how to achieve direct interaction of multimodal features in ViT and improve the performance of our tracking pipeline by using the original semantic information of the cross-modal features. Most transformer-based methods design an independent feature fusion module to preserve the powerful feature extraction and relationship modeling capabilities of ViT. Following this design, the critical problem becomes how to design a channel feature fusion module and a spatial feature fusion module.
In this work, we propose an RGB-T tracker referred to as CSTNet. CSTNet explores the advantages of directly integrating RGB and TIR features in space and channels in ViT. To the best of our knowledge, our work is the first to introduce direct interactions in ViT-based RGB-T trackers. Similar to [2], we design a cross-modal channel feature fusion module (CFM) and a spatial feature fusion (SFM) module. We cascade them into a block and insert the blocks into the ViT, which is our backbone. The CFM performs joint channel enhancement and multi-level spatial modeling of multimodal features, and then globally integrates the original semantic and fused semantic information across modal features. The SFM achieves cross-modal feature interaction through cross-attention, and then combines spatial and channel features through convolutional feedforward layers.
In addition, to promote the practicality of the model, we also propose a smaller version of CSTNet-small. It removes the CFM and SFM from the network and uses the CSTNet model as the pre-training weights. The CSTNet-small with fewer parameters and faster speed will promote the applications of our work in intelligent vehicles such as unmanned aerial vehicle (UAV), unmanned surface vehicle (USV), underwater vehicle (UV), etc.
Ours main contributions are summarized as follow:
(1) We propose a cross-modal channel feature fusion module (CFM) that first concatenates and linearly fuses RGB and TIR features, then performs channel enhancement and multi-level spatial feature modeling in parallel and sums the output features, and finally globally integrates the hybrid features with the original features.
(2) We propose a cross-modal spatial feature fusion module (SFM) that first aggregates local spatial features, then models the spatial relationships of multimodal features using a cross-attention module, and finally integrates the spatial and channel information of multimodal features through a convolutional feedforward network.
(3) We propose an RGB-T tracker CSTNet. CSTNet combines channel and spatial feature fusion with Transformer, and achieves effective cross-modal interaction by directly fusing the RGB and TIR features. Comprehensive experiments have shown that our method achieves state-of-the-art performance.
(4) We propose a smaller version CSTNet-small. Compared to CSTNet, it removes the CFM and SFM modules and uses CSTNet as the pre-training model and achieves over 36% reduction in parameters, over 24% reduction in Flops, and 50% speedup with 1-2% performance decrease.
II Related Work
II-A RGB-T tracker based on Indirect Feature Interaction
Some RGB-T trackers propose channel fusion and spatial fusion methods for cross-modal features. DAFNet [7] proposes a channel-attention-based recursive fusion chain to achieve cross-modal and multi-level feature fusion in MDNet. SiamCDA [12] proposes a complementary aware feature fusion module. It concatenates RGB and TIR features, and then enhances cross-modal features through a joint channel attention module. Finally, it achieves feature fusion through cross-modal residual connection and feature concatenation. SiamTDR [13] concatenates RGB and TIR features and then calculates the weight of channel attention by dynamic convolution to obtain fused feature. The fusion module of SiamMALL [14] is similar to the above methods, except that it replaces the feature concatenation of the last layer with feature addition. MACFT [20] proposes a cross-attention fusion module and a mix-attention fusion module for spatial feature fusion in Transformer. It only performs cross-attention operations on the multimodal features output by its backbone to obtain fused features.
Compared with these channel feature fusion modules, our method not only uses channel enhancement, but also introduces multi-level local spatial feature modeling and global integration module to achieve cross-modal feature fusion. Compared to the spatial fusion module, our method explores the potential of cross-attention and integrates the spatial representation of the fused feature by convolutional feedforward network. In addition, we insert our modules into different layers of the backbone to achieve cross-modal feature interactions at different stages. Compared to these trackers, our method achieves continuous interaction of cross modal features in channels and space, while these methods only explore channel fusion or spatial feature interaction separately.
II-B RGB-T tracker based on Indirect Feature Interaction
BAT [16] proposes a universal bi-directional adapter to achieve interaction between RGB and TIR features in transformer blocks. TBSI [2] uses the fused template as a bridge to achieve information interaction between RGB and TIR modalities in the search area based on the attention mechanism. CAT++ [21] proposes five different attribute-based branches to perceive different challenges, and then fuses the features of different branches through an aggregation interaction module to achieve multimodal tracking. TATrack [17] introduces an online template to model temporal information, and achieves cross-modal feature interaction between the initial template and the online template.
However, these methods achieve cross-modal feature fusion by using intermediate features, prior knowledge, and temporal information. They do not explore the effectiveness of direct interaction between RGB and TIR features, which results in the model not being able to benefit from the original semantic interaction between RGB and TIR features. In contrast, our method explores the potential for direct channel and spatial fusion of cross-modal features in the transformer network.
II-C Feature fusion for other RGB-T visual tasks
The channel and spatial feature fusion module has been widely applied to other non-sequential RGB-T tasks. CMX [18] uses cross-modal channel and spatial attention to complement and correct multimodal features. It also proposes a feature fusion module based on a cross-attention module and a convolutional feedforward layer for spatial modeling and joint integration of cross-modal features. MMSFormer [19] proposes a multimodal features fusion module based on multi-level feature fusion and channel enhancement. It fuses multimodal features from different stages in the backbone to obtain a robust mixed-modality feature representation of the target. SwinNet [22] performs spatial alignment based on spatial attention on cross-modal features, and then channel attention modules are used to recalibrate RGB and TIR features separately.
Compared to these methods, our method is different in terms of task form. Compared to the cross-attention module in CMX [18], our cross-attention-like module improves cross-modal feature interaction by aggregating local space features, sharing integration layers, maintaining cross-modal feature channel consistency during feature fusion, and decoupling the learning of target space and channel features.
III Method
Figure 2: The overall framework of CSTNet. Our backbone is a bilateral ViT with shared weights. The RGB and TIR image patches are first embedded into tokens and fed into 12 Transformer Blocks. The proposed CFM and SFM perform cross-modal feature fusion on template features and search area features in the network, respectively. They are inserted into 4-th, 7-th, and 10-th layers of the backbone. Finally, the RGB and TIR features of the search area are added, and the prediction head is used to predict the current state of the target.
Our method is based on the assumption of spatial alignment of RGB-T images. The overall framework of CSTNet is shown in Figure 2. We first embed the template images and search area images into tokens, and then fed the tokens into transformer blocks. Our proposed channel feature fusion module (CFM) and spatial feature fusion module (SFM) are inserted between the transformer blocks for cross-modal feature interaction. Finally, the RGB and TIR features of the search area output from the last transformer block are added and fed into the prediction head to obtain the target state. Due to the identical processing of template and search area in CFM and SFM, we only show the processing of multimodal features in the search area in sections III-B and III-C. CSTNet-small removes the CFM and SFM modules from CSTNet and trains the modal with CSTNet weights as a pre-trained model.
III-A Backbone
We follow the recent RGB-T trackers [2][17] to select ViT [23] as our backbone. The inputs of the backbone are the RGB and TIR template images, represented as $z_{r},z_{t}\in R^{H_{z}\times W_{z}\times 3}$ and the RGB and TIR search area images are represented as $x_{r},x_{t}\in R^{H_{x}\times W_{x}\times 3}$. These four images are first embedded as tokens, and positional embeds are added as follows:
$\begin{split}Z_{r}&=\text{PE}(z_{r})+E_{z},Z_{t}=\text{PE}(z_{t})+E_{z}\\ X_{r}&=\text{PE}(x_{r})+E_{x},X_{t}=\text{PE}(x_{t})+E_{x}\end{split}$ (1)
where PE is the Patch Embed function as in [23]. $Z_{r},Z_{t}\in R^{N_{z}\times C}$ denotes the init feature of RGB and TIR template tokens. $N_{z}=H_{z}W_{z}/p^{2}$, where $p$ is the down-sampling stride of the patch embedding. $C$ is the channel number. $X_{r},X_{t}\in R^{N_{x}\times C}$ denote the init feature of RGB and TIR seach area tokens, where $N_{x}=H_{x}W_{x}/p^{2}$. $E_{z}$ and $E_{x}$ are the position embeds with the same shape as $Z_{r}$ and $X_{r}$, respectively.
Since the feature extraction and modeling is the same for both modalities, we only discuss the process in the RGB modality. The RGB tokens are concatenated as $H_{r}=[X_{r};Z_{r}]\in R^{(N_{x}+N_{z})\times C}$, then $H_{r}$ is fed into a series of Transformer [23] blocks.
The proposed cross-modal feature fusion method is represented as $Z_{rt}=f(Z_{r},Z_{t}),X_{rt}=f(x_{r},x_{t})$ where $Z_{rt}$ denotes the fusion feature of the template and $X_{rt}$ denotes the fusion feature of the search area. After the fusion, the relationship modeling is represented as:
$\small\begin{split}A&=\text{Softmax}(\frac{QK^{T}}{\sqrt{C}})V=\text{Softmax}(\frac{[X_{qrt};Z_{qrt}][X_{krt};Z_{krt}]^{T}}{\sqrt{C}})V\\ &=\text{Softmax}(\frac{[X_{qrt}X_{krt}^{T},X_{qrt}Z_{krt}^{T};Z_{qrt}X_{krt}^{T},Z_{qrt}Z_{krt}^{T}]}{\sqrt{C}})V\end{split}$ (2)
where Q, K, V denotes query, key and values matrices, respectively.
From the above formulation, after the fusion, the template features and the search area features used for relationship modeling contain both RGB and TIR semantic information. The multi-level fusion of RGB and TIR features and the joint modeling of cross-modal features improve the model’s utilization of multi-level cross-modal semantic information.
III-B Channel feature fusion module
Our proposed CFM module aims to integrate RGB and TIR features of templates or search areas into the channel. As shown in Figure 3, CFM contains a channel enhancement module for jointly enhancing key information within channels, a local spatial aggregation module (LSA) for jointly modeling multi-level spatial feature representation, and a global integration module (GIM) for integrating original and fused features.
We first linearly fuse multimodal features, specifically, we concatenate the RGB and TIR features of the search area, and then use a linear layer to obtain the fused feature:
$X_{c}=\text{Linear}([X_{r};X_{t}])\in R^{N_{x}\times C}$ (3)
where Linear represents the linear layer.
Then, an squeeze-and-excitation (SE) module [24] is used as a channel enhancement unit to achieve channel enhancement of $X_{c}$.
$X_{c}^{\text{se}}=\text{SE}(X_{c})$ (4)
The local spatial aggregation module (LSA) is used to extract multi-level local spatial features and model key representations of mixed multi-level features by a non-linear layer. Specifically, a linear transformation is first performed on the feature $X_{c}$ to obtain $X_{c}^{\text{fc}}=\text{Linear}(X_{c})\in R^{N_{x}\times C}$.
Then we use convolutional blocks to extract multi-level spatial features, represented as:
$X_{c}^{\text{msf}}=\text{Conv}_{1\times 1}(X_{c}^{\text{fc}})+\sum_{k\in(3,5,7)}\text{DWConv}_{k\times k}(X_{c}^{\text{fc}})$ (5)
where Conv and DWConv represent convolutional layer with batch-normalize (BN) and depth-wise separable convolutional layer with BN, respectively.
Another linear layer is used to perform a linear transformation on the fused features to integrate multi-level features and improve their consistency, represented as $X_{c}^{\text{lsa}}=\text{Linear}(\text{Gelu}(\text{Conv}_{1\times 1}(X_{c}^{\text{msf}})))\in R^{N_{x}\times C}$, where Gelu represents the GELU activate function [25].
After the parallel SE and LSA modules, the features are added, represented as $X_{c}^{\text{add}}=X_{c}^{\text{se}}+X_{c}^{\text{lsa}}\in R^{N_{x}\times C}$. In addition, residual connections are introduced to complement the original semantic information lost in multimodal features, represented as $X_{r}^{\text{res}}=X_{r}+X_{c}^{\text{add}}$ and $X_{t}^{\text{res}}=X_{t}+X_{c}^{\text{add}}$.
The global integration module (GIM) concatenates the RGB and TIR features and feeds it into a multi-layer perceptron (MLP) with residual connection. Then, the concatenated feature is split to obtain integrated RGB and TIR features. The process is described as:
$X_{r}^{\text{cfm}},X_{t}^{\text{cfm}}=\text{Split}([X_{r}^{\text{res}};X_{t}^{\text{res}}]+\text{MLP}([X_{r}^{\text{res}};X_{t}^{\text{res}}]))$ (6)
where Split maps the feature from $R^{N_{x}\times 2C}$ to two $R^{N_{x}\times C}$ features. MLP consists of two linear layers. The $X_{r}^{\text{cfm}}$ and $X_{t}^{\text{cfm}}$ are the CFM output features.
Due to the difference between the template and the search area, GIM is not shared between the template and the search pane when processing features.
III-C Spatial feature fusion module
Our proposed SFM module aims to model spatial relationships in cross-modal features and to integrate the channel and spatial features of the RGB and TIR modalities through a convolutional feedforward network. As shown in Figure 4, the SFM contains a local perception unit (LPU), a cross-attention module, and a convolutional feedforward network (CFN).
Figure 3: Illustration of our proposed CFM module. It encompasses a channel enhancement module, which is a SE module [24], a local spatial aggregation module (LSA), and a global integration module (GIM).
Figure 4: Illustration of our proposed SFM module. It encompasses a local perception unit (LPU), a cross-attention module and a convolutional feedforward network (CFN).
We first use a local perception unit (LPU) to aggregate local spatial information, represented as:
$\begin{split}X_{r}^{\text{lpu}}=X_{r}^{\text{cfm}}+\text{DWConv}_{3\times 3}(X_{r}^{\text{cfm}})\\ X_{t}^{\text{lpu}}=X_{t}^{\text{cfm}}+\text{DWConv}_{3\times 3}(X_{t}^{\text{cfm}})\end{split}$ (7)
Then a cross-attention module is used to achieve spatial modeling of cross-modal features. Specifically, a non-linear full-connection layer is used and represented as $X_{r}^{\text{res}},X_{r}^{\text{attn}}=\text{Split}(\text{Relu}(\text{Linear}^{\text{s}}(X_{r}^{\text{lpu}})))$, where $\text{Linear}^{\text{r}}$ is a linear layer shared by RGB and TIR branches. After that, a linear layer is used to obtain Query: $X_{r}^{q}=\text{Linear}(X_{r}^{\text{attn}})$, another linear layer is used to obtain Key: $X_{r}^{k}$ and Value: $X_{r}^{v}$, represented as $X_{r}^{k},X_{r}^{v}=\text{Split}(\text{Linear}(X_{r}^{\text{attn}})$. The TIR $X_{t}^{q},X_{t}^{k},X_{t}^{v}$ are obtained through the same way.
The cross-attention process of RGB and TIR features is described as follows:
$\begin{split}X_{r}^{cross}=\text{Attention}(Q:X_{r}^{q},K:X_{t}^{k},V:X_{t}^{v})\\ X_{t}^{cross}=\text{Attention}(Q:X_{t}^{q},K:X_{r}^{k},V:X_{r}^{v})\\ \end{split}$ (8)
Then we use residual connections and a shared linear layer to adjust the features and sum the features. The reason for using sum instead of concat is to maintain channel consistency between the semantic structure of the RGB features and the TIR features of the target. The shared layer improves the learning of unified features in cross-modal features of the target. The process is described as:
$\begin{split}X_{r}^{\text{add}}&=\text{LN}_{r}(X_{r}^{\text{cfm}}+\text{Linear}^{s}(X_{r}^{cross}+X_{r}^{res}))\\ X_{t}^{\text{add}}&=\text{LN}_{t}(X_{t}^{\text{cfm}}+\text{Linear}^{s}(X_{t}^{cross}+X_{t}^{res}))\end{split}$ (9)
where $\text{LN}_{r}$ and $\text{LN}_{r}$ are different layer-normalize layers.
For the same reasons as above, we add the features $X_{r}^{\text{add}}$ and $X_{t}^{\text{add}}$ to achieve feature point-level interaction between the multimodal features.
$X_{rt}^{\text{add}}=X_{r}^{\text{add}}+X_{t}^{\text{add}}$ (10)
Due to the differences between RGB and TIR features, the directly added feature requires further spatial and channel integration. Therefore, a convolutional feedforward network (CFN) is introduced. The CFN improves the spatial and channel representation of key features by learning spatial and channel decoupling of cross-modal features. In the CFN, we first perform spatial integration on the merged feature. The process can be written as follows:
$X_{rt}^{local}=X_{rt}^{\text{add}}+\sum_{k\in(1,3)}\text{DWConv}_{k\times k}(X_{rt}^{\text{add}})$ (11)
Then, we nonlinearly integrate the channel feature of the merged feature, represented as:
$X_{rt}^{act}=X_{rt}^{local}+\text{Conv}_{1\times 1}^{\text{down}}(\text{Gelu}(\text{Conv}_{1\times 1}^{\text{up}}(X_{rt}^{local})))$ (12)
where $\text{Conv}_{1\times 1}^{\text{up}}$ maps the feature from $R^{C}$ to $R^{2\times C}$ and $\text{Conv}_{1\times 1}^{\text{down}}$ maps the feature from $R^{2\times C}$ to $R^{C}$.
In addition, we use a 1x1 convolutional block and a skip connection with a 1x1 convolutional block to adjust the features, the process is described as follows:
$X_{rt}^{adj}=\text{BN}(\text{Conv}_{1\times 1}^{\text{adj}}(X_{rt}^{act})+\text{Conv}_{1\times 1}^{\text{res}}(X_{rt}^{act}))$ (13)
where $\text{Conv}_{1\times 1}^{\text{adj}}$ and $\text{Conv}_{1\times 1}^{\text{res}}$ both map the feature from $R^{C}$ to $R^{C}$.
Finally, we add $X_{rt}^{\text{sfm}}$ to the original RGB feature $X_{r}$ and the original TIR feature $X_{t}$ to obtain the fused features as:
$X_{r}^{\text{sfm}}=X_{rt}^{adj}+X_{r},\ X_{t}^{\text{sfm}}=X_{rt}^{adj}+X_{t}$ (14)
where $X_{r}^{\text{sfm}}$ and $X_{t}^{\text{sfm}}$ are the output features of SFM.
III-D Head and Loss
The RGB and TIR features of the search area output by the last Transformer block are represented as $X_{r}^{\text{output}}$ and $X_{t}^{\text{output}}$. We add them up to obtain the mixed feature, described as:
$X_{rt}^{\text{output}}=X_{r}^{\text{output}}+X_{t}^{\text{output}}$ (15)
where $X_{rt}^{\text{output}}$ is the input of our prediction head.
Following the design of [23] and [2], we adopt a fully convolutional center-head [26] to predict the state of the target.
We use the weight focal loss [27] to train the classification branch and the l1 loss and the wise-IoU loss [28] to train the box regression branch. The total training loss of CSTNet is:
$L_{total}=L_{cls}+\lambda_{iou}L_{iou}+\lambda_{l_{1}}L_{1}$ (16)
where $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are hyper-parameters as same in [2].
IV Experiments
TABLE I: Comparison with state-of-the-art trackers on RGBT210 [29], RGBT234 [30] and LasHeR [31] testing set. Higher value Indicate better performance. The best two results are shown in red and blue fonts.
Method Source Baseline RGBT210 [29] RGBT234 [30] LasHeR [31]
PR SR PR SR PR NPR SR
TFNet [32] TCSVT2021 MDNet 77.7 52.9 80.6 56.0 - - -
DMCNet [33] TNNLS22 MDNet 79.7 55.9 83.9 59.3 49.0 43.1 35.5
APFNet [9] AAAI2022 MDNet 79.9 54.9 82.7 57.9 50.0 43.9 36.2
X-Net [34] Arxiv2023 MDNet - - 85.2 62.2 50.8 - 44.6
CAT++ [21] TIP2024 MDNet 82.2 56.1 84.0 59.2 50.9 44.4 35.6
SiamCDA [12] TCSVT2021 Siamese - - 76.0 56.9 - - -
DFAT [35] Inf Fusion23 Siamese - - 75.8 55.2 - - -
SiamMLAA [14] TMM2023 Siamese - - 78.6 58.4 - - -
MFNet [36] IVC2022 DCF - - 84.4 60.1 59.7 55.4 46.7
LSAR [37] TCSVT2023 DCF - - 78.4 55.9 - - -
CMD [38] CVPR2023 DCF - - 82.4 58.4 59.0 54.6 46.4
AMNet [6] TCSVT2024 DCF - - 85.5 60.7 61.4 55.9 47.7
ProTrack [15] ACM MM2022 Transformer 79.5 59.9 78.6 58.7 53.8 - 42.0
ViPT [1] CVPR2023 Transformer 83.5 61.7 83.5 61.7 65.1 - 52.5
MACFT [20] Sensors2023 Transformer - - 85.7 62.2 65.3 - 52.5
RSFNet [39] ISPL2023 Transformer - - 86.3 62.2 65.9 - 52.6
DFMTNet [40] IEEE Sens.J23 Transformer - - 86.2 63.6 65.1 - 52.0
TBSI [2] CVPR2023 Transformer 85.3 62.5 87.1 63.7 69.2 65.7 55.6
STMT [41] Arxiv2024 Transformer 83.0 59.5 86.5 63.8 67.4 63.4 53.7
TATrack [17] Arxiv2024 Transformer 85.3 61.8 87.2 64.4 70.2 66.7 56.1
BAT [16] AAAI2024 Transformer - - 86.8 64.1 70.2 - 56.3
TransAM [6] TCSVT2024 Transformer - - 87.7 65.5 70.2 66.0 55.9
CSTNet-small - Transformer 85.1 62.1 86.7 64.1 70.3 66.3 56.4
CSTNet - Transformer 86.0 63.5 88.4 65.2 71.5 67.9 57.2
IV-A Datasets and Evaluation Metrics
LasHeR [31] is a large-scale dataset benchmark that contains 1224 sequences with 730K frame pairs. It is divided into 19 challenging attributes, which are No Occlusion (NO), Partial Occlusion (PO), Total Occlusion (TO), Hyaline Occlusion (HO), Out-of-View (OV), Low Illumination (LI), High Illumination (HI), Abrupt Illumination Variation (AIV), Low Resolution (LR), Deformation (DEF), Background Clutter (BC), Similar Appearance (SA), Thermal Crossover (TC), Motion Blur (MB), Camera Moving (CM), Frame Lost (FM), Fast Motion (FM), Scale Variation (SV), Aspect Ratio Change (ARC). RGBT234 [30] is a classic benchmark for RGB-T tracking that contains 234 sequences. RGBT234 [30] has 12 challenging attributes, they are no occlusion (NO), partial occlusion (PO), heavy occlusion (HO), low illumination (LI), low resolution (LR), thermal crossover (TC), deformation (DEF), fast motion (FM), scale variation (SV), motion blur (MB), camera moving (CM), and background clutter (BC). RGBT210 [29] is another classic RGB-T tracking benchmark that contains 210 sequences. VTUAV [42] is a large-scale RGB-T tracking dataset in aerial view and its short-term tracking test subset contains 176 sequences. GTOT [43] is a classic RGB-T tracking benchmark consisting of 50 short-term sequences.
Following the metrics used by current most RGB-T trackers, precision rate (PR) and success rate (SR) are employed to evaluate different trackers on RGBT210 [29], RGBT234 [30], and GTOT [43], and precision rate (PR), normalized precision rate (NPR) and success rate (SR) are employed to evaluate the trackers on LasHeR [31], and maximum precision rate (MPR) and maximum success rate (MSR) are employed to evaluate the trackers on VTUAV [42].
Figure 5: The SR and PR scores of CSTNet, TBSI [2] and BAT [16] under different attributes on LasHeR [31].
Figure 6: The SR and PR scores of CSTNet, TBSI [2], mfDiMP [44], under different attributes on RGBT234 [30].
IV-B Implementation Details
CSTNet is implemented on Ubuntu 20.04 using Python 3.9 and Pytorch 1.13.0. It is trained on a platform with two NVIDIA RTX A6000 GPUs over 20 epochs, which took roughly 6 hours. The LasHeR [31] training set is used to train our model. In each epoch, we sample 60k image pairs, with a total batch-size of 64. The total learning rate is set to $2e^{-5}$ for the backbone and $2e^{-6}$ for feature fusion modules. It decays 10× after 15 epochs. We use AdamW as the optimizer with $1e^{-4}$ weight decay. The sizes of template and search area are set to 128×128 and 256×256, respectively. Our CFM and SFM modules are cascaded and inserted into the 4-th, 7-th, and 10-th layers of the ViT.
CSTNet-small removes the CFM and SFM modules from the network and uses the CSTNet weights as the pre-training weights, while its training settings are the same as CSTNet.
IV-C Comparison with State-of-the-art Methods
IV-C1 Comparison methods
To evaluate the effectiveness of our method, we compare CSTNet and CSTNet-small with RGB-T trackers from recent years, including 5 MDNet-based trackers, including TFNet [32], DMCNet [33], APFNet [9], X-Net [34], CAT++ [21], 3 Siamese-based trackers, including SiamCDA [12], SiamMALL [14], DFAT [35], 4 DCF-based trackers MFNet [36], LSAR [37], CMD [38], AMNet [6], and 11 transformer-based trackers, including ProTrack [15], ViPT [1], MACFT [20], RSFNet [39], DFMTNet [40], TBSI [2], STMT [41], BAT [16], TATrack [17], TransAM [6].
TABLE II: Comparison with state-of-the-art trackers on GTOT [43]. The best two results are shown in red and blue fonts.
Baseline Backbone GTOT
PR SR
TFNet [32] MDNet VGG-M 72.4 88.6
APFNet [9] MDNet VGG-M 73.7 90.5
DMCNet [33] MDNet VGG-M 73.8 90.9
SiamCDA [12] Siamese Res-50 87.7 73.2
DFAT [35] Siamese Res-50 89.3 72.3
SiamMLAA [14] Siamese Res-50 91.3 75.1
LSAR [37] DCF VGG-16 70.3 85.5
CMD [38] DCF Res-18 89.2 73.4
AMNet [6] DCF Res-50 91.1 76.1
RSFNet [39] Transformer ViT-B 92.1 75.3
TransAM [6] Transformer ViT-B 92.9 77.4
CSTNet-small Transformer ViT-B 91.5 75.4
CSTNet Transformer ViT-B 93.0 76.7
TABLE III: Comparison of parameters, Flops, and FPS on GPU GTX 3090Ti and GPU GTX 2080.
3090Ti / 2080 Params Flops
torch onnxruntime
TBSI [2] 60 / 28 71 / 40 201.98M 82.5G
CSTNet 62 / 36 74 / 44 144.00M 74.7G
CSTNet-small 94 / 47 126 / 65 92.12M 56.4G
Figure 7: Comparison with state-of-the-art trackers on VTUAV [42]. The trackers include HMFT [42], Stark [45], TransT [46], mfDiMP [44], FSRPN [47], TCTrack [48], ADRNet [49], DAFNet [7].
IV-C2 Evaluation on LasHeR Dataset
The trackers are evaluated on LasHeR [31] and the results are reported in Table I. CSTNet achieves state-of-the-art performance. Compared to other Transformer-based trackers, CSTNet outperforms TATrack [17] and TransAM [6] by 1.3% in PR, 1.2% and 1.9% in NPR, and 1.1% and 1.3% in SR, respectively. In addition, CSTNet’s PR and SR are 1.3% and 0.9% higher than BAT [16]. In addition, CSTNet surpasses all other MDNet-based and Siamese-based methods. CSTNet-small achieves performance close to that of BAT. The results show that CSTNet achieves better cross-modal feature interaction and better performance.
In addition, we evaluate CSTNet, TBSI [2], and BAT [16] on 19 attributes of the LasHeR [31] dataset. As shown in Figure 5, CSTNet outperforms other transformer-based trackers on many attributes. Compared to TBSI [2], CSTNet achieves more competitive performance in challenges such as occlusion (HO, PO, and SO), thermal crossover (TC), out-of-view (OV) and motion blur (MB). Compared to BAT [16], CSTNet has better performance in attributes such as heavy occlusion (HO), abrupt illumination variation (AIV), and background clutter (BC). However, the PR and SR of CSTNet on ARV, FL, SA, AIV, and LI attributes are insufficient, which proves that CSTNet needs to further improve the robustness and feature discrimination of the model.
IV-C3 Evaluation on RGBT234 Dataset
The trackers are evaluated on RGBT234 [30] and the results are reported in Table I. Compared to the TransAM [6], CSTNet improves the PR by 0.7%. Compared to the BAT [16], CSTNet improves the PR and SR by 1.6% and 1.1%, respectively. CSTNet outperforms CSTNet-small by 1.7% in PR and 1.1% in SR. The results show that CSTNet sets a new state-of-the-art benchmark.
We further evaluate our proposed CSTNet, TBSI [2], and mfDiMP [44] on 12 attributes of the RGBT234 [30] dataset. As shown in Figure 6, CSTNet achieves better performance on most attributes compared to other trackers. However, it has an insufficient success rate in heavy occlusion (HO), and low resolution (LR) background clutter (BC), and fast motion (FM). This indicates that it has shortcomings in dealing with challenges such as missing key features and interference from background features.
IV-C4 Evaluation on RGBT210 Dataset
The trackers are evaluated on RGBT210 [29] and the results are reported in Table I. CSTNet outperforms all other RGB-T trackers. The PR, SR of CSTNet are 0.7%, 1.0% higher than TBSI [2]. In addition, CSTNet achieves an SR of 63.5%, outperforming STMT [41] and TATrack [17] by 4.0% and 1.7%, respectively. The experimental results demonstrate that CSTNet outperforms all other trackers and achieves state-of-the-art performance.
IV-C5 Evaluation on VTUAV-ST Dataset
The trackers are evaluated on VTUAV [42] and the results are reported in Table 7. CSTNet and CSTNet-small outperform all other RGB-T trackers. Compared to HMFT [42], CSTNet improves the MPR by 1.7% and the MSR by 3.3%. Compared to CSTNet, CSTNet-small improves the MPR by 4.7% and the MSR by 3.3%. The experimental results demonstrate that CSTNet-small outperforms all other trackers and achieves state-of-the-art performance.
IV-C6 Evaluation on GTOT Dataset
The trackers are evaluated on GTOT [43] and the results are reported in Table II. CSTNet outperforms all other RGB-T trackers in PR. Compared to RSFNet [39], CSTNet improves the PR and SR by 0.9% and 1.4%, respectively. The experimental results demonstrate that CSTNet and CSTNet-small achieves state-of-the-art performance.
TABLE IV: Ablation studies of our proposed CFM and SFM modules in the tracking pipeline.
Method PR NPR SR
RGBT 70.0 66.3 56.0
RGBT+CFM 70.5 (+0.5) 67.0 (+0.7) 56.5 (+0.5)
RGBT+SFM 68.5 (-1.5) 64.8 (-1.5) 55.1 (-0.9)
RGBT+CFM&SFM 71.5 (+1.5) 67.9 (+1.6) 57.2 (+1.2)
TABLE V: Ablation of the different methods for fusing search area cross-modal features.
Method PR NPR SR
concat 70.0 66.4 55.1
add 71.5 67.9 57.2
TABLE VI: Ablation of the different components of CFM.
CFM PR NPR SR
SE LSA GIM
69.5 65.8 55.8
70.5 66.8 56.3
70.8 67.1 56.7
71.5 67.9 57.2
TABLE VII: Ablation of the different components of SFM.
SFM PR NPR SR
LPU Cross-Attn CFN
70.5 67.0 56.5
70.6 67.1 56.5
71.0 67.4 56.8
71.5 67.9 57.2
TABLE VIII: Ablation studies of different insert layers of our CFM and SFM modules.
Layers PR NPR SR
4 7 10
70.0 66.3 56.0
70.4 66.6 56.3
70.9 67.1 56.8
71.5 67.9 57.2
TABLE IX: Ablation studies of different pre-trained models.
Pre-train Model PR NP SR
- 55.6 51.7 42.2
ImageNet [50] 59.1 56.0 47.3
SOT [23] 69.7 65.9 55.8
RGBT [2] 71.5 67.9 57.2
IV-D Parameters and Flops
We also evaluate the Parameters, Flops, and FPS of the models. The results are shown in Table III. Compared to our baseline [2], our CSTNet and CSTNet-small achieve better performance, reducing approximately 25% and 50% of the parameters and 10% and 30% of the Flops, respectively. Our CSTNet and CSTNet-small achieve 62 FPS and 94 FPS, respectively, on the GPU GTX 3090Ti. We use onnxruntime to further accelerate the model. The accelerated CSTNet-small achieves a speed of 126 FPS. We also test the speed of the trackers on the GPU GTX2080. CSTNet and CSTNet-small run at a speed of 36 and 47 FPS respectively. After the deployment, their speeds reached 44 FPS and 65 FPS respectively. In addition, we deploy our CSTNet and CSTNet-small on NVIDIA Jetson Xavier. CSTNet and CSTNet-small run at speeds of 21 FPS and 33 FPS, demonstrating the feasibility of our method in practical applications.
IV-E Ablation Studies
To investigate the individual influence of different components, we conduct ablation studies on CSTNet and report the results for PR, NPR and SR on LasHeR [31].
IV-E1 Component Analysis
First, we evaluate the effect of summing the RGB and TIR features and the CFM and SFM modules on the performance of the tracker. RGBT denotes adding the RGB and TIR features of the search area output by the shared ViT backbone, and then predicting the target state. The pretrain-model of the baseline is [2]. As shown in Table IV, an improvement of 0.5% and 1.0% in terms of PR, 0.7% and 0.9% in terms of NPR, and 0.5% and 0.7% in terms of SR is obtained by the contribution of the CFM module and the SFM module, respectively. In addition, they achieve a total improvement of 1.5%, 1.6%, and 1.2% in terms of PR, NPR, and SR, respectively. In addition, the 1-2% performance degradation of CSTNet when using only the SFM module confirms the discussion in TBSI that cross-attention may not be suitable for RGB-T tracking. In fact, the cross-modal features integrated through CFM can lead to positive gains with the cross-attention based SFM module. This further illustrates the importance of cascading CFM and SFM modules. They are interdependent, rather than simply cascading two modules to improve features separately, which distinguishes our method from other cross-attention-based methods. Overall, both our proposed CFM and SFM modules are important.
In addition, we separately investigate the impact of concatenating or adding the RGB and TIR features of the search area output from the backbone on the tracking performance. concat is used by TBSI [2], which means that the RGB and TIR features of the search area are concatenated and then a 1x1 convolution is used to reduce the dimension. As shown in Table V, the addition of RGB and TIR features in the search area achieves better performance. Our method avoids the feature representation bottleneck caused by the 1x1 convolution dimension reduction, allowing the model to benefit from the direct addition of RGB and TIR features in the search area.
IV-E2 Ablation of CFM
We evaluate the contributions of each component of the CFM. As shown in Table VI, when the CFM module is removed from the model, the PR and SR are 69.5% and 55.8%, respectively. Adding the SE module improves the PR by 1.0% and the SR by 0.5%. Then, using the LSA module improves performance by 0.3% for SR and 0.4% for PR. Next, using the GIM module to integrate global features results in an improvement of 0.7% for PR and 0.5% for SR. Overall, all three modules in CFM contribute to improving the performance of our model.
IV-E3 Ablation of SFM
The contribution of each component to the SFM is measured. Table VII shows the results. The contribution of the LPU is a 0.1% improvement in PR. The cross-attention module improves the PR by 0.4% and SR by 0.3%. The CFN module improves the PR by 0.5% and SR by 0.4%. Overall, all three components of the SFM are effective.
Figure 8: Visualization of heat maps of RGB feature, TIR feature and fusion feature of the search area on LasHeR [31] based on Grad-CAM. The Fusion Heat Map represents the heatmap of the sum feature of the RGB and TIR modalities. The fusion feature’s heat map is only displayed in RGB modality. (a) bikeboyright (frame 301), (b) biketurnright (frame 46), (c) whitebikebelow (frame 165), (c) yellowgirl118 (frame 65).
Figure 9: Visualization of heat maps of the fusion feature of CSTNet, TBSI [2] and BAT [16] on LasHeR [31] based on Grad-CAM. The fusion feature’s heat map is only displayed in RGB modality. (a) basketboy (frame 48), (b) bikefromlight (frame 138), (c) motocomeonlight (frame 31), (c) blkstandboy (frame 128).
IV-E4 Inserting Layers of CFM and SFM
In addition, we conduct ablation experiments on different insert layers of our proposed CFM and SFM modules. The results are reported in Table VIII. Inserting the CFM and SFM modules in the 4th layer of the ViT improves the PR by 0.4% and the SR by 0.3%. Inserting the modules in the 7th and 10th layers of ViT further improves the performance of the tracker. Overall, inserting our CFM and SFM modules into these three layers effectively achieves feature fusion at different stages and improves model performance.
IV-E5 Ablation of pre-trained models
We explore the impact of different pre-trained models on our proposed CSTNet in terms of performance. CSTNet without pretraining and CSTNet with ImageNet pretraining are unable to converge in 20 epochs, therefore we increase their learning rate by 10 times. The results are shown in Table IX, CSTNet with RGB-T pretraining achieves the best performance. It outperforms CSTNet with SOT pretraining by 1.8% for PR and 1.4% for SR. The RGB-T pre-training model achieves better parameter initialization of our model, reducing our training epochs to one-third compared to ViPT [1], TATrack [17], and STMT [41] and BAT [16]. This shows that using the RGB-T pre-trained model can further improve the performance of the tracker.
The use of pre-trained RGB-T models is not intended to gain an inappropriate performance advantage; rather, we hope to promote the development of RGB-T trackers through a better pre-trained model, as in the case of [23].
IV-F Visualization
IV-F1 Heatmap
We explore how CSTN utilizes RGB and TIR modalities to improve model performance. As shown in Figure 8, in the bikeboyright sequence, which is a typical daytime traffic scenario, our proposed CSTNet effectively uses RGB and TIR modalities to better predict the target state. In the biketurnright sequence which is a typical nighttime scenario, although CSTNet is disturbed in RGB feature, the model still makes a correct prediction in fused features under the reference of TIR feature. In the whitebikebelow and yellowgirl118 sequences, CSTNet improves target state prediction in high illumination and partial occlusion challenges by complementing the features of RGB and TIR modalities. CSTNet effectively achieves cross-modal feature interaction.
In addition, we compare the cross-modal feature interaction performance of CSTNet and other Transformer-based methods (TBSI [2] and BAT [16]) on four typical sequences of the LasHeR [31] dataset, as shown in Figure 9. Our proposed CSTNet achieves direct interaction between RGB and TIR original semantics through channel and spatial fusion of multimodal features. In the basketboy, bikefromlight, and motocomeonlight sequences, CSTNet achieves richer utilization of cross-modal appearance features. In the blkstandboy sequence, CSTNet has better discrimination for similar appearance. Overall, CSTNet demonstrates the benefits of direct cross-modal feature interaction on model performance.
IV-F2 Tracking Results
Figure 10: Visualized comparisons of CSTNet with TBSI [2], BAT [16] and APFNet [9] on four sequences from LasHeR [31] dataset.
We select four representative sequences from the Lasher dataset to show some of the results of CSTNet and other previous state-of-the-art RGB-T trackers. As shown in Figure 10, in the rightdarksingleman sequence, when other trackers briefly lose the target due to low light and camera shake, our proposed CSTNet still tracks the target stably, demonstrating the better robustness of CSTNet. In the ab_blkskirtgirl sequence, when the target is frequently occluded, other trackers usually lose the target briefly, while our CSTNet does not lose the target. In the bikefromlight sequence, when the scene is highly illuminated, our CSTNet accurately distinguishes the target from the surrounding clutter, resulting in better performance. In the boytakingplate2left sequence, our CSTNet better integrates the RGB and TIR features of the target, resulting in better performance in partial occlusion and similar appearance challenges. Overall, our proposed CSTNet has stronger discriminative ability and better tracking performance.
V Conclusion
In this paper, we present a new RGB-T tracker CSTNet. CSTNet explores the advantages of using channel and spatial feature fusion to achieve direct interaction between RGB and TIR features. CSTNet contains two novel cross-modal feature interaction modules, CFM and SFM. The CFM achieves joint channel enhancement of RGB and TIR features, joint multi-level modeling, and global integration of fused and original features. The SFM achieves spatial mutual modeling of RGB and TIR features and joint spatial and channel integration of multimodal features. Comprehensive experiments show that the direct interaction between RGB and TIR features improves the performance of the model, and our proposed CSTNet achieves state-of-the-art performance on three challenging RGB-T tracking benchmarks. We also provide a small version, CSTNet-small. It achieves high performance RGB-T tracking using the model without the CFM and SFM modules and CSTNet weights as the pre-training weights. We evaluate the parameters, flops, and speed of CSTNet and CSTNet-small, and deploy them to a embedded computing device. Comprehensive experiments demonstrate that our models meet the real-time requirement of practical applications.
VI Acknowledgments
This research is funded by the National Key Research and Development Program of China, grant number 2023YFC2809104, by the National Natural Science Foundation of China, grant number 52371350, and by the National Key Laboratory Foundation of Autonomous Marine Vehicle Technology, grant number 2024-HYHXQ-WDZC03.